Say you have a sphere within a cube. The height of the cube is 12 cm and the circumference of the sphere is 7 cm. Find the volume of the area within the cube that is not within the sphere. Round your answer to the nearest tenth.
Answer by jorel1380(3719)

The cube's area is 12x12x12cc's=1728 cm3.
The circumference of the sphere, 2πr=7, so r=7/2π=1.115cms.
The volume of the sphere is 4/3π(1.115)^3=5.806cm3
The empty area around the sphere is 1722.2 cm3..